What do you know about 2 dice being rolled?
You know that the sample space of one die being rolled is
{1,2,3,4,5,6}, if you roll two dice,
you have a sample space with 36 elements. Here it is: (1,1)(1,2)(1,3)(1,4)(1,5)(1,6)(2,1)(2,2)(2,3)(2,4)(2,5)(2,6)(3,1)(3,2)(3,3)(3,4)(3,5)(3,6)(4,1)(4,2)(4,3)(4,4)(4,5)(4,6)(5,1)(5,2)(5,3)(5,4)(5,5)(5,6)(6,1)(6,2)(6,3)(6,4)(6,5)(6,6)
From this we can calculate many probabilities. We could look at the sums. For example what is the probability of rolling two dice and have a sum of 2?
There is only one way out of 36 for this to happen so it is 1/36.
Similarly, what is the chance of rolling a 1 and a 3, in either order?
What is the chance of having a sum of 6? All these answers come directly from
the sample space above.
Name Three possible outcomes of obesity?
Diabities, blood clot, being made fun of, embarrassment, appearance ect.
When rolling a number cube once what is the probability of rolling a 3 in percentage?
(1/6)*100%= 16.67%
z = 0.84 approx.
What is the standard deviation of a sample of 35 25 36 42 37?
Putting the data into excel, the std dev for the sample is 6.20
How many distinct three letter arrangements can be found from the letters in mathematics?
The number of different three letter arrangements that can be done from the
letters in the word "mathematics"
is; 11P3 =
11!/(11-3)! =
990
Keeping score of a set of observations?
you could use a table saying when you saw it on 1 side and where on the other side
Does Non-parametric statistics fall in a normal distribution that reflect an ordinal scale?
Not necessarily.
.05*.95*.95=0.045125=4.5125%
What is the probability of rolling two dice that the first dice is an even number?
There are only three even numbers on a "legal" six sided dice (2,4 & 6), which ever dice is designated as "first" whether rolled independently or in unison will have a 1 in 2 chance of being even.
Put another way: legal dice, designate first, 3 even, 3 odd, 50/50.
Bonus Answer: Applies to second dice also.
Compute the probability given an event that a male person will get pregnant mathematically?
The probability that mathematics will make a male pregnant is zero!
What is the estimd population mean of sample 36 mean weight 3.01 standard diviation 03?
The answer will depend on the population mean of what variable? Height?, length or is it simply weight. If it is weight, the estimated (not estimd) population mean is 3.01 units: the same as the sample mean. The standard deviation (not diviation) is irrelevant.
The answer will depend on the population mean of what variable? Height?, length or is it simply weight. If it is weight, the estimated (not estimd) population mean is 3.01 units: the same as the sample mean. The standard deviation (not diviation) is irrelevant.
The answer will depend on the population mean of what variable? Height?, length or is it simply weight. If it is weight, the estimated (not estimd) population mean is 3.01 units: the same as the sample mean. The standard deviation (not diviation) is irrelevant.
The answer will depend on the population mean of what variable? Height?, length or is it simply weight. If it is weight, the estimated (not estimd) population mean is 3.01 units: the same as the sample mean. The standard deviation (not diviation) is irrelevant.
What is the best way to study probability in math?
It is best to study it systematically and carefully. Make sure you understand the various concepts before moving on to more complicated ideas.
It is best to study it systematically and carefully. Make sure you understand the various concepts before moving on to more complicated ideas.
It is best to study it systematically and carefully. Make sure you understand the various concepts before moving on to more complicated ideas.
It is best to study it systematically and carefully. Make sure you understand the various concepts before moving on to more complicated ideas.
What is the probability that they will have at least one girl if a couple plans to have 6 children?
There is no simple answer to the question because the children's genders are not independent events. They depend on the parents' ages and their genes.
However, if you assume that they are independent events then, given that the probability of a girl is approx 0.48, then
Prob(at least one girl out of 6 children) = 1 - Prob(no girls out of 6 children)
= 1 - Prob(6 boys out of 6 children)
= 1 - 0.526 = 0.9809
There is no simple answer to the question because the children's genders are not independent events. They depend on the parents' ages and their genes.
However, if you assume that they are independent events then, given that the probability of a girl is approx 0.48, then
Prob(at least one girl out of 6 children) = 1 - Prob(no girls out of 6 children)
= 1 - Prob(6 boys out of 6 children)
= 1 - 0.526 = 0.9809
There is no simple answer to the question because the children's genders are not independent events. They depend on the parents' ages and their genes.
However, if you assume that they are independent events then, given that the probability of a girl is approx 0.48, then
Prob(at least one girl out of 6 children) = 1 - Prob(no girls out of 6 children)
= 1 - Prob(6 boys out of 6 children)
= 1 - 0.526 = 0.9809
There is no simple answer to the question because the children's genders are not independent events. They depend on the parents' ages and their genes.
However, if you assume that they are independent events then, given that the probability of a girl is approx 0.48, then
Prob(at least one girl out of 6 children) = 1 - Prob(no girls out of 6 children)
= 1 - Prob(6 boys out of 6 children)
= 1 - 0.526 = 0.9809
What is the function of the chromatain?
The chromatain have four major functions. They package DNA into a smaller volume to fit in the cell. They strengthen the DNA to allow mitosis, and they prevent damage to DNA. Chromatain control gene expression and DNA replication.
Odds.
The probability of getting 5 on one toss of a pair is 1/9. The most probable number of fives you will get after 42 tosses is 4 or 5.
Add the probabilities of getting 5 four, five, six, seven, eight etc times out of 42 tosses. By the time you get to 15 times, the probability will be very small and you can quit adding them up.
The probability of rolling 5 four times is (1/9)^4*(8/9)^38*42!/(38!*4!)
= 119,930*1.524*10^-4*1.138*10^-2
= 0.20803
The probability of rolling 5 five times is (1/9)^5*(8/9)^37*42!/(37!*5!)
= 8.507*10^5*1.6935*10^-5*1.2804*10^-2
= 0.18447
I am not going to continue this because it would take half the night, but you get the idea.
There is an easier way to do this that just occurred to me. It will give the same answer. Here it is:
Add the probabilities of getting 5 zero, one, two or three times, and subtract that from 1.
The probability of never getting five is (8/9)^42 = 0.007105
The probability of getting 5 only once is (1/9)*(8/9)^41*42 = 0.037304
The probability of getting 5 twice is
(1/9)^2*(8/9)^40*(42*41*/2) = 0.095591
Now YOU compute the probability of getting 5 three times. Add that to the probabilities above and subtract the sum from one. You should end up with a number around 0.6.
That will be the answer.
What is an important difference between the uniform and normal probability distributions?
The uniform distribution is limited to a finite domain, the normal is not.
What are the odds of rolling a four on one die?
Since there are six sides on a die, the odds of rolling a 4, or any digit for that matter is 1/6
What is the function of the common bileduct?
To drain bile from the liver (and the gall bladder) into the small intestines.
(Bile is a combination of waste products (mostly from decomposing hemoglobin) and solvents designed to help dissolve (digest) fats in your food.)
What are the chances of becoming a freelance writer in California?
Freelancing is a home based job, you can do it from any where and serve customers of any location.
What you need to have is excellent writing skills and analytical skills. If you are looking for working as a freelancer, you can choose Obuhi and get hired as per your skills they will help you to find a good job.